OpenStudy (anonymous):

Find the volume of the solid obtained by rotating the region R bounded by the curves y = x and about the x-axis.

OpenStudy (amistre64):

y=x and ???

OpenStudy (anonymous):

oops! y = x and y = 2sqrtx

OpenStudy (anonymous):

something missing... y=x & what else? what is the boundaries? x=[0,?]

OpenStudy (anonymous):

got it! give me sec

OpenStudy (amistre64):

that better :) pi [S] x^2 - 4x dx from some interval lol

OpenStudy (amistre64):

x^3/3 - 2x^2

OpenStudy (anonymous):

x=[0,4]

OpenStudy (amistre64):

x = 2sqrt(x) x^2 = 4x x^2 - 4x = 0 x = 4/2 +- sqrt(16)/2 x = 2 +- 2; so o to 4

OpenStudy (amistre64):

maaa!! iniks reading my mind again ;)

OpenStudy (anonymous):

lol!

OpenStudy (amistre64):

64/3 - 32 is the answer

OpenStudy (anonymous):

you are good with this one - I'll fish for something else. enjoy!

OpenStudy (amistre64):

96 64 --- 32/3 = 10' 2/3

OpenStudy (amistre64):

32/3 is the answer if i did it right ;)

OpenStudy (anonymous):

Amistre HELP!!!

OpenStudy (amistre64):

where you at loco :)

OpenStudy (anonymous):

the third question i just posted it

OpenStudy (anonymous):

thanks amistre!

OpenStudy (anonymous):

ur welcome :)